Show that Tn=X+1/n+1 is a consistent estimator of the parameter theta of a binomial population


1
Expert's answer
2021-06-02T12:49:03-0400

"T_n=\frac{x+1}{n+1}"

An estimator "T_n"  of parameter θ is consistent if it converges in probability:

"\displaystyle{\lim_{n\to \infin}}Pr(|T_n-\theta|>\varepsilon)=0"

for all "\varepsilon>0"

For a binomial population: "\theta=x/n" - the proportion of successes.

Then:

"\displaystyle{\lim_{n\to \infin}}(T_n-\theta)=\displaystyle{\lim_{n\to \infin}}(\frac{x+1}{n+1}-\frac{x}{n})=0"


If

"(\frac{x+1}{n+1}-\frac{x}{n})=0"

then:

"Pr(|\frac{x+1}{n+1}-\frac{x}{n}|>\varepsilon)=0"

because "\varepsilon>0" .

So:

"\displaystyle{\lim_{n\to \infin}}Pr(|\frac{x+1}{n+1}-\frac{x}{n}|>\varepsilon)=0"